CONICAL INTERSECTIONS IN THE PRESENCE OF SPIN-ORBIT COUPLING

\underline{SPIRIDOULA MATSIKA}, DAVID R. YARKONY, Department of Chemistry, Johns Hopkins University, Baltimore MD 21218.

Conical intersections do not exist as isolated points but as continuous seams of dimension N^{int} − 2 for the nonrelativistic Coulomb hamiltonian, where N^{int} is the number of internal coordinates. Including the spin-orbit interaction can change the situation qualitatively. For odd-electron systems the hamiltonian is complex causing the dimension to reduce to N^{int} − 3 or even N^{int} − 5 when there is no spatial symmetry present. In order to be able to study the effect of the spinorbit interaction quantitatively we have implemented an algorithm that can locate conical intersections in the presence of the spin-orbit interaction. The system H_{2} + OH has been found to have conical intersections between ^{2}Σ−^{2}Π states in linear symmetry or 1^{2}A′−2^{2}A′ states in C_{s} symmetry, in the nonrelativistic case. Points of the seam including the spin-orbit coupling have been located and compared with the nonrelativistic results.